OpenStudy (anonymous):

The average speed of one airplane is 200 mph faster than that of an automobile. If the automobile starts at 3 a.m. and the airplane starts from the same place at 11 a.m., they will have traveled the same number of miles by 1:00 p.m. Find the rate of the automobile. i got 25 mph?

OpenStudy (anonymous):

I am getting 50 mph.

OpenStudy (saifoo.khan):

\[\Huge \color\red {Congrats!}\]

OpenStudy (saifoo.khan):

:)

OpenStudy (anonymous):

Let x = rate of automobile Let x + 200 = rate of plane time automobile travels = 3am to 1pm = 10 hours time plane travels = 11am to 1pm = 2 hours rate of auto x time auto travels = rate of plane x time plane travels x(10) = (x + 200)(2)

OpenStudy (anonymous):

10x = 2x + 400 8x = 400 x = 50

OpenStudy (anonymous):

You probably didn't multiply the 200 by 2. :)